OpenStudy (anonymous):

A right triangle has two legs with lengths a and b and a hypotenuse with leg c. In this triangle, the area and perimeter are the same nonzero number. Find the length of a if b=6. (Hint: use the Pythagorean Theorum and solve algebraically).

myininaya (myininaya):

of you know the area of a triangle and the perimeter write those formulas given you know b=6

myininaya (myininaya):

we will also need the Pythagorean theorem

myininaya (myininaya):

what you need: A=.5ab P=a+b+c a^2+b^2=c^2 you are given b=6 You are also given A=P so let's rewrite those: P=.5a(6)=3a P=a+(6)+c Use these two equations to write an equation in terms of two variables. Then use that equation and sub into a^2+6^2=c^2

myininaya (myininaya):

let me know if you still don't understand

OpenStudy (anonymous):

hi myininaya, i think i understand, but i'm confused on how to combine the two equations. c introduces another variable which i believe complicates things. thank you so much for your help!! :)

myininaya (myininaya):

So A and P are equal that is why we can set this equation 3a=a+c+6 2a=c+6 so c=2a-6 now go to a^2+6^2=c^2 and sub the c in this equation and solve for a

myininaya (myininaya):

you will only have one variable

myininaya (myininaya):

So you will find a then you go back to c=2a-6 to find c

OpenStudy (anonymous):

so i plug in 2a-6 in for c in the pythagorean theorum?

myininaya (myininaya):

:) yep

OpenStudy (anonymous):

thank you so much! oh wait whad do i plug in for a ????

myininaya (myininaya):

you are solving that equation for a

OpenStudy (anonymous):

and does it plug in for c or c^2?

myininaya (myininaya):

nothing to plug in for a you are trying to come up with an equation that contains one variable for c just replace the c part make sure you keep that exponent 2 because we haven't done anything with it yet

OpenStudy (anonymous):

do i also square the six?

myininaya (myininaya):

we have a^2+6^2=c^2 replacing c with 2a-6 with a^2+6^2=(2a-6)^2 You need to multiply (2a-6)(2a-6)

OpenStudy (anonymous):

ah! thank you bunches, if i could give you a million medals i would, i think i finally understand!!!!!!

myininaya (myininaya):

good news :)

OpenStudy (anonymous):

so i got it to a=4a^2-24a. i was never very good at algebra, do i put the 24a on the other side and then find the square root of both sides, and then divide both sides by the square root of the right side?

myininaya (myininaya):

is that a^2=4a^2-24a?

OpenStudy (anonymous):

yes. i just do not know what to do with it from here

myininaya (myininaya):

Right it as a degree 2 polynomial =0

myininaya (myininaya):

Or get everything on one side is what i'm saying

OpenStudy (anonymous):

so now i have 25a-4a^2=0???

myininaya (myininaya):

I thought you had a^2=4a^2-24a?

OpenStudy (anonymous):

yes i did, but then i moved everything to one side

myininaya (myininaya):

so a^2-4a^2+24a=0

myininaya (myininaya):

a^2 and 24a aren't liek terms

OpenStudy (anonymous):

OH I FORGOT TO SQUARE THE A

myininaya (myininaya):

a^2 and -4a^2 are

OpenStudy (anonymous):

but that would give you -3a -4a^2

myininaya (myininaya):

no wouldn't...

myininaya (myininaya):

what is a^2-4a^2?

OpenStudy (anonymous):

oh it would give you -3a^2+24a

myininaya (myininaya):

yes -3a^2+24a=0

myininaya (myininaya):

now both of the terms have -3a in common so factor -3a out

OpenStudy (anonymous):

i got a^2-8a

OpenStudy (anonymous):

do i leave it like that?

OpenStudy (anonymous):

OH IS A=8!

myininaya (myininaya):

so so if you factored out -3a why do you have a^2 there?

myininaya (myininaya):

and yes a=8

myininaya (myininaya):

-3a^2+24a=0 3a(a-8)=0 a-8=0 a=8

OpenStudy (anonymous):

thank you you are brilliant :)))))))))

myininaya (myininaya):

because obviously one of the lengths can't be the other answer which is a=0

myininaya (myininaya):

it wouldn't be a triangle lol

OpenStudy (anonymous):

i did it differently but still got the same answer?

OpenStudy (anonymous):

i divided both sides by just -3, and got a^2-8a=0, and then i moved the 8a to the other side to make it positive, and then divided both sides by a

myininaya (myininaya):

well just so you know usually you can't divide both sides by a because a can be zero but a actually cannot be 0 here so that is fine just put next to that step a not equal to 0